OpenStudy (anonymous):

Find the perimeter of a rectangular object which has a length of √(75) feet and a width of √(48) feet

OpenStudy (anonymous):

\[18\sqrt{3}\]

OpenStudy (anonymous):

Can you please explain, in complete sentences, how you arrived at the answer and give the final solution in simplified radical form.

OpenStudy (anonymous):

The perimeter of a rectangle is 2a+2b where a and b are the sides So 2\[2\sqrt{75}+2\sqrt{48}=2\sqrt{3*25}+2\sqrt{3*16}\] \[=2*5*\sqrt{3}+2*4*\sqrt{3}=18\sqrt{3}\]

OpenStudy (anonymous):

What items around the house would be this ?

OpenStudy (anonymous):

Andras come back! haha

OpenStudy (anonymous):

For this it is easier to think of \[\sqrt{75}=8.66 and \sqrt{48}=6.93\]

OpenStudy (anonymous):

I am European so I count in meters and cm. 8.66feet= 264cm 6.93feet=211 cm That is a bed I guess

OpenStudy (anonymous):

double bed

OpenStudy (anonymous):

or a big cupboard